A broadside array consisting of 200 cm wavelength with 10 half-wave dipole spacing 10 cm. And if each array element feeding with 1 amp. current and operating at same frequency then find the half power beamwidth


1
Expert's answer
2021-01-05T12:18:32-0500

Answer

half power beam width is

HPBW=BWFN/2="\frac{2\lambda}{nd2}=\frac{\lambda}{nd}"

According to question

"\frac{\lambda}{d}=20"


"HPBW=\frac{20}{10}=2^°"